OpenStudy (anonymous):

Which of the following is a solution to 5√x - 7 = 10 ?

hero (hero):

5√(x - 7) = 10 √(x - 7) = 10/5 √(x - 7) = 2 x - 7 = 4 x = 4 + 7 x = 11

hero (hero):

I'm pretty sure my solution is correct.

OpenStudy (anonymous):

5√x - 7 = 10 5√x = 10 + 7 = 17 5√x = 17 √x = 17/5 = 3.4 x = (3.4)square = 11.56 i hope ur question syntax is fine. do tell me when you have the answer.

hero (hero):

sigh

hero (hero):

x = 11 is the solution. Original poster just forgot to put parentheses

OpenStudy (anonymous):

If you plug in 29/4, you get 2.5, not 10..

OpenStudy (anonymous):

The OP used \[5\sqrt{x-7}=10\] Not \[5\sqrt{x-5}=10\]

hero (hero):

The original poster awarded me with the medal by the way.

OpenStudy (anonymous):

@Luis Besides, even with your equation, the solution is x=9. @Hero, okay, I just wanted to make sure the the OP had the correct answer. :)